OpenStudy (anonymous):

A temperature of 41 degrees F es equivalent to 278.2 degrees K, while a temperature of 86 degrees F is equivalent to 303.2 degrees K. Find the equation converting Kelvin temperature to Fahrenheit. What is the Fahrenheit temperature when it is 180 degrees K? HELP!!! I'm stressing out!

hartnn (hartnn):

this relation is linear,so it can be written in the form of y=mx+c where y is temperature in Fahrenheit and x is temperature in Kelvin so now u have 2 points (278.2,41) and (303.2,86) can u find the eqaution using these 2 points??

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

no luck

OpenStudy (anonymous):

Sorry I'm so stupid. I can't see the equation to solve. Please don't leave me stranded.

OpenStudy (anonymous):

what is m and what is c?

hartnn (hartnn):

m is slope or proportionality constant here,c is y-intercept or offset here

OpenStudy (anonymous):

I'm sorry, I just can't figure out the equation. I hate Algebra. It hates me back.

hartnn (hartnn):

so when i put those two points,i get 2 simultaneous equation and then after solving them,i get m and c values.... so lets put point (278.2,41) in the equation y=mx+c 41=278.2*m+c -----> (1) for point (303.2,86) 86= 303.2*m+c------->(2) solving these two simultaneously,we get m=1.8 (or 9/5) and c=-459.76 so your conversion equation is F=9/5*K-459.76

OpenStudy (anonymous):

I love you and you are my hero. I am a fan now and how do I give you a medal? You are soooooooo nice for helping out an old grandma!

hartnn (hartnn):

now if u put K=180(given) in that equation,u get temperature in F as F=9/5*180-459.76=324-459.76=-135.76 degree F

OpenStudy (anonymous):

I wish you were my professor. :(